OpenStudy (anonymous):

a) Given that cos x = sqrt3 - 1 find the value of cos 2x in the form a+b sqrt3 where a and b are integers b)Given that 2 cos(y+30) = sqrt 3 sin(y-30) find the value of k sqrt3

OpenStudy (anonymous):

a) \[ x = \sqrt{3} - 1\] b) \[ 2cos(y+30) = \sqrt{3} sin (y-30)\]

OpenStudy (anonymous):

find tan in the form \[k \sqrt{3}\]

OpenStudy (anonymous):

cos2x = cos^2 x-sin^2 x = 2 cos^2 x - 1 Cosx=root3 -1, cos^2 x= 3-2root3+1 = 4-2root3, 2cos^2 x -1 = 7-4root3

OpenStudy (anonymous):

I'd answer the other one but I have to leave.